Based on the amounts of taxable income​ provided, compute the federal income tax payable in 2018 on each amount assuming the taxpayers are married filing a joint return.​ Also, for each amount of taxable​ income, compute the average tax rate and the marginal tax rate.

a. taxable income of $72,000

b. taxable income of $90,000

c. taxable income of $380,000

d. taxable income of $690,000

This is all the information that was given.
Answer #1

First let me tell you the Income Tax brackets for married, jointly filing (for taxes due in 2018) based on the Taxable Income (TI); are as follows

Upto $19050 = 10% of TI

$19051 - $77400 = $1905 + 12% of TI in excess of $19050

$77401 - $165000 = $8907 + 22% of TI in excess of $77400

$165001 - $315000 = $28179 + 24% of TI in excess of $165000

$315001 - $400000 = $64179 + 32% of TI in excess of $315000

$400001 - $600000 = $91379 + 35% of TI in excess of $400000

$600001 or more = $161379 + 37% of TI in excess of $600000

Now, based on the above tax brackets, we can compute the tax liability for all the cases. Also, for the average tax, we need to divide the total tax by the total income; and for Marginal tax rate, we need to divide the incremental tax by the incremental income.

A. Taxable Income = $72000

Tax Payable = $1905 + 12% of ($72000 - $19050) = $8259

Average Rate of Tax = $8259 / $72000 = 11.47%

Marginal Rate of Tax = 11.47% (same as average rate if tax in this case)

B. Taxable Income = $90000

Tax Payable = $8907 + 22% of ($90000 - $77400) = $11679

Average Rate of Tax = $11679 / $90000 = 12.98%

Marginal Rate of Tax = ($11679 - $8259) / ($90000 - $72000) = 19%

C. Taxable Income = $380000

Tax Payable = $64179 + 32% of ($380000 - $315000) = $84979

Average Rate of Tax = $84979 / $380000 = 22.36%

Marginal Rate of Tax = ($84979 - $11679) / ($380000 - $90000) = 25.28%

D. Taxable Income = $690000

Tax Payable = $161379 + 37% of ($690000 - $600000) = $194679

Average Rate of Tax = $194679 / $690000 = 28.21%

Marginal Rate of Tax = ($194679 - $84979) / ($690000 - $380000) = 35.39%